### v2

@@ 3.3 Multi-Modality Transformer
The MMT adopts an encoder-decoder structure, as shown in Fig. 2(b). To better model sequence contexts and capture dependency between different windows in clip, we use Transformer Vaswani et al. (2017) decoder layers to form the backbone of the decoder instead of using CNN layers. Notably, we also design intra-cross and inter-cross attention layers, respectively, to handle information interaction within and between modalities.
@@
We reshape $x_{0}$ to $(N,128)$ after the diffusion process, and the Transformer models the contextual relationships between multiple windows in a clip to ensure temporal consistency in the generated reactions. These labels include the 3DMM coefficients, denoted as $3dmm\in{\mathbb{R}}^{N\times 58}$, and emotion, denoted as ${emo}\in{\mathbb{R}}^{N\times 25}$. The basic idea behind 3DMM is to represent a 3D face as a linear combination of basic shapes and textures learned from a dataset of 3D scans. To visualize the generated facial reactions, FaceVerseV2 Wang et al. (2022) and PIRender Ren et al. (2021) are utilized, enabling the conversion of 3DMM coefficients into 2D facial images that are conditioned on the reference portrait identity. The emotion includes three attributes: VA is a 2-dimensional vector ranging from -1 to 1, AU is a 15-dimensional binary vector of 0 or 1, and FE is an 8-dimensional vector ranging from 0 to 1.
@@ 3.4 Latent Diffusion
This noise prediction network $\epsilon_{\theta}$ is a modified version of UNet following Preechakul et al. Preechakul et al. (2022). Training is done by optimizing $L_{eps}$ with respect to $\theta$:
@@ 4.1 Datasets
We have used two datasets with video-conference setting provided by REACT2024 challenge Song et al. (2024).
  • NoXI. The Novice eXpert Interaction (NoXI) dataset Cafaro et al. (2017) is a dyadic interaction dataset that is annotated during an information retrieval task targeting multiple languages, multiple topics, and the occurrence of unexpected situations. NoXI is a corpus of screen-mediated face-to-face interactions recorded at three locations (France, Germany and UK), spoken in seven languages (English, French, German, Spanish, Indonesian, Arabic and Italian) discussing a wide range of topics.
  • RECOLA. The REmote COLlaborative and Affective (RECOLA) Ringeval et al. (2013) dataset consists of 9.5 hours of audio, visual, and physiological recordings of online dyadic interactions between 46 French speaking participants.
@@ 4.2 Implementation Detail
We have trained MMT and ReactDiff on an NVIDIA A100 GPU with a batch size of 16 and learning rate of 5e-3, and 100 epochs both, respectively. Video frame resolution is $224\times 224$ and training time of two stages is 16h. We use the AdamW optimizer Loshchilov and Hutter (2016) with $\beta_{1}=0.9,\beta_{2}=0.999$ and apply a weight decay of $45e-$ 4. We generate $\alpha$ = 10 reaction sequences per speaker sequence for evaluation, evaluation metrics shown in the Appendix.
@@
Trans-VAE Song et al. (2023) 0.03 92.81
BeLFusion (k=1) Barquero et al. (2023) 0.10 92.32
BeLFusion (k=10) Barquero et al. (2023) 0.12 91.60
REGNN Xu et al. (2023) 0.19 84.54
VQ-Diff Nguyen et al. (2024) 0.30 91.65
Unifarn Liang et al. (2023) 0.19 98.51
Beamer Hoque et al. (2023) 0.11 97.33
FRDiff Yu et al. (2023) 0.14 91.05
PerFRDiff Zhu et al. (2024) 0.38 94.72
Liu et.al Liu et al. (2024) 0.22 100.43
@@ 4.3 Comprehensive Evaluation of ReactDiff’s Performance
To demonstrate the performance of the ReactDiff framework, we conducted experiments on two sub-challenges of REACT2024 Song et al. (2024): Offline and Online Multiple Appropriate Facial Reaction Generation (Offline and Online MAFRG). Offline MAFRG task focuses on generating multiple appropriate facial reaction videos from the input speaker behaviour. Online MAFRG task focuses on the continuous generation of facial reaction frames based on current and previous speaker behaviour. Some reactions generated by offline ReactDiff are shown in Fig. 4. The challenge establishes a set of naive baselines, namely B_Random, B_Mime, and B_MeanSeq/B_MeanFr. Specifically, B_Random randomly samples $\alpha$ = 10 facial reaction sequences from a Gaussian distribution. B_Mime generates facial reactions by mimicking the corresponding speaker’s facial expressions. For B_MeanSeq and B_MeanFr, the generated facial reactions are decided by the sequence-wise and frame-wise average reaction in the training set, respectively. Additionally, there are Trans-VAE Song et al. (2023), BeLFusion Barquero et al. (2023) and REGNN Xu et al. (2023) baselines. The visualization results of ReactDiff and the above methods are shown in Fig. 3.
@@ 4.3.2 Diversity result
Trans-VAE Song et al. (2023) 0.0008 0.0002 0.0006
BeLFusion (k=1) Barquero et al. (2023) 0.0068 0.0073 0.0094
BeLFusion (k=10) Barquero et al. (2023) 0.0105 0.0082 0.0116
REGNN Xu et al. (2023) 0.0007 0.0061 0.0342
VQ-Diff Nguyen et al. (2024) 0.0743 0.0348 0.0745
Unifarn Liang et al. (2023) 0.0819 0.0760 0.0259
Beamer Hoque et al. (2023) 0.0508 0.0374 0.0196
FRDiff Yu et al. (2023) 0.0690 0.0443 0.0108
PerFRDiff Zhu et al. (2024) 0.1368 0.2191 0.0879
Liu et.al Liu et al. (2024) 0.1675 0.0535 0.1385
@@ 4.3.3 Realism and synchrony result
Trans-VAE Song et al. (2023) 67.74 43.75
BeLFusion (k=1) Barquero et al. (2023) 79.60 44.94
BeLFusion (k=10) Barquero et al. (2023) 79.59 44.87
REGNN Xu et al. (2023) - 41.35
VQ-Diff Nguyen et al. (2024) - 45.33
Unifarn Liang et al. (2023) - 46.11
Beamer Hoque et al. (2023) - 48.12
FRDiff Yu et al. (2023) 69.37 47.66
PerFRDiff Zhu et al. (2024) 47.62 45.28
Liu et.al Liu et al. (2024) - 44.54
@@ Appendix A Evaluation Metrics
    where $N$ is frame number of clip, $CCC$ denotes the Concordance Correlation Coefficient Lawrence and Lin (1989), $y_{i}^{n}$ denotes the $n$-th frame of the $i$-th generated listener facial reaction ($emotion$ in paper, includes multi-channel facial attributes), $\hat{y}^{n}$ denotes the $n$-th frame of a set of appropriate real listener facial reactions corresponding to a speaker behaviour (The similarity matrix is computed for each speaker reaction sequence, and the listener reaction corresponding to multiple speakers whose similarity with the $i$-th speaker exceeds the threshold is utilized as a set of appropriate facial listener reactions corresponding to the $i$-th speaker).
  • Facial Reaction Distance (FRDist). we compute the Dynamic Time Warping (DTW) Berndt and Clifford (1994) distance between the generated facial reaction and its most similar appropriate real facial reaction, and final distance score for evaluation is obtained by averaging as:
@@
  • Facial reaction realism (FRRea). The realism score is obtained by computing Frechet Inception Distance (FID) Heusel et al. (2017) between the distribution Dis($Y$) of the generated facial reactions and the distribution Dis($\hat{Y}$) of the corresponding appropriate real facial reactions as:
@@
  • Synchrony between generated facial reactions and speaker behaviours (FRSyn). we compute the Time Lagged Cross Correlation (TLCC) Shen (2015) to evaluate the synchrony between the input speaker behaviour and the corresponding generated facial reaction.

### v3

@@ 3.3 Multi-Modality Transformer
The MMT adopts an encoder-decoder structure, as shown in Fig. 2(b). To better model sequence contexts and capture dependency between different windows in clip, we use Transformer (Vaswani et al., 2017) decoder layers to form the backbone of the decoder instead of using CNN layers. Notably, we also design intra-cross and inter-cross attention layers, respectively, to handle information interaction within and between modalities.
@@
We reshape $x_{0}$ to $(N,128)$ after the diffusion process, and the Transformer models the contextual relationships between multiple windows in a clip to ensure temporal consistency in the generated reactions. These labels include the 3DMM coefficients, denoted as $3dmm\in{\mathbb{R}}^{N\times 58}$, and emotion, denoted as ${emo}\in{\mathbb{R}}^{N\times 25}$. The basic idea behind 3DMM is to represent a 3D face as a linear combination of basic shapes and textures learned from a dataset of 3D scans. To visualize the generated facial reactions, FaceVerseV2 (Wang et al., 2022) and PIRender (Ren et al., 2021) are utilized, enabling the conversion of 3DMM coefficients into 2D facial images that are conditioned on the reference portrait identity. The emotion includes three attributes: VA is a 2-dimensional vector ranging from -1 to 1, AU is a 15-dimensional binary vector of 0 or 1, and FE is an 8-dimensional vector ranging from 0 to 1.
@@ 3.4 Latent Diffusion
This noise prediction network $\epsilon_{\theta}$ is a modified version of UNet following Preechakul et al. (Preechakul et al., 2022). Training is done by optimizing $L_{eps}$ with respect to $\theta$:
@@ 4.1 Datasets
We have used two datasets with video-conference setting provided by REACT2024 challenge (Song et al., 2024).
  • NoXI. The Novice eXpert Interaction (NoXI) dataset (Cafaro et al., 2017) is a dyadic interaction dataset that is annotated during an information retrieval task targeting multiple languages, multiple topics, and the occurrence of unexpected situations. NoXI is a corpus of screen-mediated face-to-face interactions recorded at three locations (France, Germany and UK), spoken in seven languages (English, French, German, Spanish, Indonesian, Arabic and Italian) discussing a wide range of topics.
  • RECOLA. The REmote COLlaborative and Affective (RECOLA) (Ringeval et al., 2013) dataset consists of 9.5 hours of audio, visual, and physiological recordings of online dyadic interactions between 46 French speaking participants.
@@ 4.2 Implementation Detail
We have trained MMT and ReactDiff on an NVIDIA A100 GPU with a batch size of 16 and learning rate of 5e-3, and 100 epochs both, respectively. Video frame resolution is $224\times 224$ and training time of two stages is 16h. We use the AdamW optimizer (Loshchilov and Hutter, 2016) with $\beta_{1}=0.9,\beta_{2}=0.999$ and apply a weight decay of $45e-$ 4. We generate $\alpha$ = 10 reaction sequences per speaker sequence for evaluation, evaluation metrics shown in the Appendix.
@@
Trans-VAE (Song et al., 2023) 0.03 92.81
BeLFusion (k=1) (Barquero et al., 2023) 0.10 92.32
BeLFusion (k=10) (Barquero et al., 2023) 0.12 91.60
REGNN (Xu et al., 2023) 0.19 84.54
VQ-Diff (Nguyen et al., 2024) 0.30 91.65
Unifarn (Liang et al., 2023) 0.19 98.51
Beamer (Hoque et al., 2023) 0.11 97.33
FRDiff (Yu et al., 2023) 0.14 91.05
PerFRDiff (Zhu et al., 2024) 0.38 94.72
Liu et.al (Liu et al., 2024) 0.22 100.43
@@ 4.3 Comprehensive Evaluation of ReactDiff’s Performance
To demonstrate the performance of the ReactDiff framework, we conducted experiments on two sub-challenges of REACT2024 (Song et al., 2024): Offline and Online Multiple Appropriate Facial Reaction Generation (Offline and Online MAFRG). Offline MAFRG task focuses on generating multiple appropriate facial reaction videos from the input speaker behaviour. Online MAFRG task focuses on the continuous generation of facial reaction frames based on current and previous speaker behaviour. Some reactions generated by offline ReactDiff are shown in Fig. 4. The challenge establishes a set of naive baselines, namely B_Random, B_Mime, and B_MeanSeq/B_MeanFr. Specifically, B_Random randomly samples $\alpha$ = 10 facial reaction sequences from a Gaussian distribution. B_Mime generates facial reactions by mimicking the corresponding speaker’s facial expressions. For B_MeanSeq and B_MeanFr, the generated facial reactions are decided by the sequence-wise and frame-wise average reaction in the training set, respectively. Additionally, there are Trans-VAE (Song et al., 2023), BeLFusion (Barquero et al., 2023) and REGNN (Xu et al., 2023) baselines. The visualization results of ReactDiff and the above methods are shown in Fig. 3.
@@ 4.3.2 Diversity result
Trans-VAE (Song et al., 2023) 0.0008 0.0002 0.0006
BeLFusion (k=1) (Barquero et al., 2023) 0.0068 0.0073 0.0094
BeLFusion (k=10) (Barquero et al., 2023) 0.0105 0.0082 0.0116
REGNN (Xu et al., 2023) 0.0007 0.0061 0.0342
VQ-Diff (Nguyen et al., 2024) 0.0743 0.0348 0.0745
Unifarn (Liang et al., 2023) 0.0819 0.0760 0.0259
Beamer (Hoque et al., 2023) 0.0508 0.0374 0.0196
FRDiff (Yu et al., 2023) 0.0690 0.0443 0.0108
PerFRDiff (Zhu et al., 2024) 0.1368 0.2191 0.0879
Liu et.al (Liu et al., 2024) 0.1675 0.0535 0.1385
@@ 4.3.3 Realism and synchrony result
Trans-VAE (Song et al., 2023) 67.74 43.75
BeLFusion (k=1) (Barquero et al., 2023) 79.60 44.94
BeLFusion (k=10) (Barquero et al., 2023) 79.59 44.87
REGNN (Xu et al., 2023) - 41.35
VQ-Diff (Nguyen et al., 2024) - 45.33
Unifarn (Liang et al., 2023) - 46.11
Beamer (Hoque et al., 2023) - 48.12
FRDiff (Yu et al., 2023) 69.37 47.66
PerFRDiff (Zhu et al., 2024) 47.62 45.28
Liu et.al (Liu et al., 2024) - 44.54
@@ Appendix A Evaluation Metrics
    where $N$ is frame number of clip, $CCC$ denotes the Concordance Correlation Coefficient (Lawrence and Lin, 1989), $y_{i}^{n}$ denotes the $n$-th frame of the $i$-th generated listener facial reaction ($emotion$ in paper, includes multi-channel facial attributes), $\hat{y}^{n}$ denotes the $n$-th frame of a set of appropriate real listener facial reactions corresponding to a speaker behaviour (The similarity matrix is computed for each speaker reaction sequence, and the listener reaction corresponding to multiple speakers whose similarity with the $i$-th speaker exceeds the threshold is utilized as a set of appropriate facial listener reactions corresponding to the $i$-th speaker).
  • Facial Reaction Distance (FRDist). we compute the Dynamic Time Warping (DTW) (Berndt and Clifford, 1994) distance between the generated facial reaction and its most similar appropriate real facial reaction, and final distance score for evaluation is obtained by averaging as:
@@
  • Facial reaction realism (FRRea). The realism score is obtained by computing Frechet Inception Distance (FID) (Heusel et al., 2017) between the distribution Dis($Y$) of the generated facial reactions and the distribution Dis($\hat{Y}$) of the corresponding appropriate real facial reactions as:
@@
  • Synchrony between generated facial reactions and speaker behaviours (FRSyn). we compute the Time Lagged Cross Correlation (TLCC) (Shen, 2015) to evaluate the synchrony between the input speaker behaviour and the corresponding generated facial reaction.